A girl running at a constant speed of 1.4m/s in a straight line throw a ball directly upward at 4.7 m/s. How far will she travel before the ball drops back into her hand? Ignore air fiction.

Answer by ankor@dixie-net.com(12701)

A girl running at a constant speed of 1.4m/s in a straight line throw a ball directly upward at 4.7 m/s. How far will she travel before the ball drops back into her hand? Ignore air fiction.
:
First find out how long the ball will be in the air,
t = seconds
Gravitational constant: -4.9 m/sec^2
Upward velocity: +4.7 m/s
:
-4.9t^2 + 4.7t = 0
t(-4.9t + 4.7) = 0
Solutions:
t = 0
and
-4.9t = -4.7
t = -4.7/-4.9
t = +.959 seconds
:
Find out how far he runs in .959 sec at a rate of 1.4 m/s
:
.959 * 1.4 = 1.346 m
:
He travels 1.346 meters from the time the ball is thrown until it is caught